OpenStudy (steve816):

School tomorrow and I still haven't finished my summer math packet. PLEASE DON'T MAKE MY MATH TEACHER KILL ME!!! Fogell owns a classic car whose value has been increasing by 8.5% each year. The car is worth $5000. He wants to sell the car when it is worth $10,000. How many years to the nearest hundredth must he wait?

OpenStudy (anonymous):

\[10000=5000\left( 1+\frac{ 8.5 }{ 100 } \right)^n\] \[\frac{ 10000 }{ 5000 }=\left( \frac{ 1085 }{ 1000 } \right)^n\] \[\left( \frac{ 217 }{ 200 } \right)^n=2\] find n

OpenStudy (steve816):

Wow, brilliant!

OpenStudy (anonymous):

This is beyond me I have no idea, surjithayer sounds about right.